OpenStudy (anonymous):

I tried solving for K but i got the wrong answer so can someone help please and thank you... solve for K 0.337=(15.2+k) over 48. i got 0.337 but it was wrong ???

OpenStudy (anonymous):

First of all @Nelhaj1 : A Warm Welcome to 'Open Study'. Please Read CoC (compulsory to read by all "Open Study" users) : http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

.337 = (15.2 + k)/48 multiply the equation by 48 16.176 = 15.2 + k subtract 15.2 from each side .976 = k

OpenStudy (anonymous):

do you understand what I did ?

OpenStudy (anonymous):

the anaswer is 0.976

OpenStudy (anonymous):

Yes I did redid it using your method just so i could understand it thank you very much:)

OpenStudy (anonymous):

thats what we are here for :)

OpenStudy (petiteme):

0.337=(15.2+k) over 48 0.337 * 48= 15.2 + k (0.337 * 48) - 15.2 = k :)

OpenStudy (anonymous):

thank you @Blaze @kelliegirl33 @Dakota005 @petiteme

OpenStudy (anonymous):

anytime :)

OpenStudy (anonymous):

:)

OpenStudy (petiteme):

you're welcome! :) anytime :)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

never leave openstudy

OpenStudy (anonymous):

Your Welcome Proud to Welcome a New comer at anytime